OpenStudy (bruno102):

Write an equation of a line in slope intercept form that is perpendicular to the line 2x -3y = 12 and passes through the point (2, 6). Having a lot of trouble with this.

OpenStudy (anonymous):

is this on connections?

OpenStudy (anonymous):

first, set your equation in slope interrupt form.... do you know how to ?

OpenStudy (anonymous):

intercept *

OpenStudy (bruno102):

No it is not @sWeetesttea

OpenStudy (bruno102):

Do you mean y=mx+b?

OpenStudy (bruno102):

@megzie

OpenStudy (anonymous):

yes

OpenStudy (bruno102):

Ok, so now what do I do?

OpenStudy (anonymous):

what did you get?

OpenStudy (bruno102):

I do not know how to turn this equation into slope intercept form

OpenStudy (anonymous):

ok so right now you have: 2x-3y=12 and slope intercept form is: y=mx+b This is what each letter means: 'm' is slope, this is a number you get with the 'x' symbol that is next to it. 'b' is the y-intercept, this is the number without a variable next to it. Looking at the slope intercept form We can see that the 'y' is alone on the left. That's what we need to try and do. we have: 2x-3y=12, what do you do to get the '-3y' alone...?

OpenStudy (bruno102):

Divide by -3?

OpenStudy (bruno102):

@megzie

OpenStudy (anonymous):

no, This is our equation: 2x-3y=12 you have a '2x' on the same side as the 'y', what do you do with the '2x' so that the '-3y' can be alone?

OpenStudy (bruno102):

So subtract 2x from both sides. -3y=12-2x

OpenStudy (anonymous):

The answer is y=(-3/2)x + 9 :)

OpenStudy (bruno102):

Thanks! But how did you get it?